# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 3 Extension to the Unbounded Case
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}},2(C+1)\right)\vee\log 2,$
then $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
To summarize, if condition (3.5) holds for $\varepsilon_{J}$, then we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\frac{a_{J}}{1-a_{J}}$
$\leq 6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Thus, (3.5) implies a lower bound on $P(A_{J}^{c})$.
Now, set $\kappa=\frac{3(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}}$. Suppose $J$ satisfies condition (3.5). Then $J\leq J^{\ast}$ with $f^{\ast}$ the output of $J^{\ast}$ steps of the algorithm. By the Cauchy sequence argument in (2.2), we have $\|f^{\ast}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-2}}$. Now if we assume further that $\|\Upsilon_{K}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-1}}$, i.e., the event $A_{J}^{c}$, we have by the triangle inequality that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{3d}{2^{J-1}}=\kappa\varepsilon_{J}$. If we call $B_{J}$ the event that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\kappa\varepsilon_{J}$, we have $P(A_{J}^{c})\leq P(B_{J})$ for $J$ satisfying (3.5). We already produced a lower bound on $P(A_{J}^{c})$ for such $J$. So in fact, for all $J$ satisfying (3.5), we have a lower bound on $P(B_{J})$, namely, that
$\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\kappa\cdot\varepsilon_{J}$
with probability at least
$1-6\cdot\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Let $J^{\ast}$ be the maximal $J$ for which conditions (3.5) all hold, with $J^{\ast}=1$ if none do. Note that $\varepsilon_{J}\downarrow 0$ as $J\to\infty$, so (3.5) will not hold for all sufficiently large $J$. This implies $J^{\ast}<\infty$. Further, set $\underline{C}=6$ and $C^{\prime}=\frac{1}{2}$. Then for any $1\leq J\leq J^{\ast}$ including $J^{\ast}=1$, we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J}\right)$ $\leq\underline{C}\cdot\mathbbm{1}(J>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$ (3.6)
Note the first inequality trivially holds for $J=1$ since $\kappa\varepsilon_{1}=3d$, so that the probability will be 0 as will the right-hand side. Similarly, we extend this to hold for all $J\in\mathbb{Z}$ with $J\leq J^{\ast}$, as if $J\leq 0$, then $\kappa\cdot\varepsilon_{J}=\frac{3d}{2^{J-1}}\geq 6d$. But since $d$ is the diameter of $\mathcal{F}$, the above probability is 00 which is less than than the right hand side of (3.6).
Now observe that $\bigcup_{J\leq J^{\ast}}[\varepsilon_{J},\varepsilon_{J-1})=[\varepsilon_{J^{\ast}},\infty)$. Pick any $x\geq\varepsilon_{J^{\ast}}$. Then $x\in[\varepsilon_{J},\varepsilon_{J-1})$ for some $J\leq J^{\ast}$. Then
$2\kappa x\geq 2\kappa\cdot\varepsilon_{J}=2\kappa\cdot\varepsilon_{J-1}.$
Note further that $x\mapsto\exp(-C^{\prime}nx^{2})$ is non-increasing for $x\geq 0$. Consequently,
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J-1}\right)$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J-1}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2}).$
Thus $\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2})$ for all $x\geq\varepsilon_{J^{\ast}}$.
Then we integrate:
$\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $=\int_{0}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$
$\leq C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+\int_{2\kappa\varepsilon_{J^{\ast}}}^{\infty}2x\cdot\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime\prime}nx^{2})\mathrm{d}x$
$=C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+C^{\prime\prime\prime\prime}\cdot\mathbbm{1}(J^{\ast}>1)n^{-1}\exp(-C^{\prime\prime\prime\prime\prime}n{\varepsilon_{J^{\ast}}}^{2}).$
Recall from our assumption (3.5) that either $J^{\ast}=1$ or $n{\varepsilon_{J^{\ast}}}^{2}>\log 2$. In either case, $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on $\sigma$, $d$, and our constant $C>3$.
The following theorem then brings together the lower and upper bounds we obtained.
Theorem 3.6.
Suppose $\mathcal{F}$ satisfies (3.1). Define $\varepsilon^{\ast}:=\sup\{\varepsilon:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\}$ where $c$ is a sufficiently large absolute constant. Then the minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to absolute constants.
{toappendix}
Proof of Theorem 3.6.
The proof is entirely identical to Theorem 2.12, except we replace $L(c/2-1,\sigma,B_{\mathcal{F}})$ from the bounded case with $\widetilde{L}(c/2-1,\sigma)$ everywhere. E.g., we set $\eta=\frac{c\sqrt{2}}{\widetilde{L}(c/2-1,\sigma)}\wedge 1\in(0,1)$. ∎
4 Adaptive Version
We relax the definition of local metric entropy by removing the supremum and instead considering the pointwise value of the local metric entropies. This lets us produce an estimator that is adaptive to the true point $\bar{f}$, which is a difference from the work in yang1999information and mendelson_2017_local_versus_global, and to our knowledge is the first example of an adaptive estimator in this general setting.
Definition 4.1 (Adaptive Local Entropy).
Let $M(r,\mathcal{F})$ be the largest cardinality of a $r$-packing set in the $L_{2}(\mathbb{P}_{X})$ norm of $\mathcal{F}$. Let $c>0$. Then we define
$M_{\mathcal{F}}^{\operatorname{adloc}}(f,\varepsilon,c)=M(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F}).$
Then $\log M^{\operatorname{adloc}}_{\mathcal{F}}(f,\varepsilon,c)$ is the adaptive local entropy of $\mathcal{F}$ at $f$.
A following lemma relates the adaptive local metric entropy of two nearby points.
Lemma 4.2.
Let $f,g\in\mathcal{F}$ satisfy $\|f-g\|_{L_{2}(\mathbb{P}_{X})}<\delta$. Then $M_{\mathcal{F}}^{\operatorname{adloc}}(f,\varepsilon,c)\leq M_{\mathcal{F}}^{\operatorname{adloc}}(g,2\varepsilon,2c)$ provided $\varepsilon>\delta$.
Proof.
If we show $B(f,\varepsilon)\subseteq B(g,2\varepsilon)$, then the cardinality of the largest $(\varepsilon/c)$-packing of $B(f,\varepsilon)\cap\mathcal{F}$ will be less than or equal to the cardinality of the largest $(2\varepsilon/2c)$-packing of $B(g,2\varepsilon)\cap\mathcal{F}$. Well, if $h\in B(f,\varepsilon)$, then $\|h-f\|_{L_{2}(\mathbb{P}_{X})}<\varepsilon$. Using the triangle inequality,
$\|h-g\|_{L_{2}(\mathbb{P}_{X})}\leq\|h-f\|_{L_{2}(\mathbb{P}_{X})}+\|g-h\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon+\delta<2\varepsilon,$
showing that $h\in B(g,2\varepsilon)$. ∎
We now show that the main estimator is adaptive to the true point $\bar{f}$, in which the key condition (4.2) does not require taking a supremum over all $f\in\mathcal{F}$ in the metric entropy term. The proof works no matter if $\mathcal{F}$ is bounded or not in sup-norm and also applies in the sub-Gaussian noise setting.
Theorem 4.3.
Consider the bounded case. Let $\overline{J}$ be the maximal integer $J\geq 1$ with $\varepsilon_{J}=\frac{d\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}}{2^{J-2}c}$ satisfying
@@ 4 Adaptive Version
$n\varepsilon_{J}^{2}>\inf_{f\in\mathcal{F}}2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(f,\frac{2\varepsilon_{J}\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},2c\right)\vee\log 2.$ (4.1)
and $\overline{J}=1$ if this condition is never satisfied. Let $f^{\ast}(\vec{X},\vec{Y})$ be the output of $\overline{J}$ steps of Algorithm 1.
@@
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{2\varepsilon_{J}\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},2c\right)\vee\log 2,$ (4.2)
noting the use of the true $\bar{f}\in\mathcal{F}$. Then $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on universal constants, $B_{\mathcal{F}}$, and $\sigma$. We defined $L(C,\sigma,B_{\mathcal{F}})$ in (2.3), where $C$ is some sufficiently large universal constant and $c:=2(C+1)$ is the constant we use in the definition of local metric entropy.
In the unbounded case, assuming $\mathcal{F}$ satisfies (3.1), simply replace every instance of $L(C,\sigma,B_{\mathcal{F}})$ with $\widetilde{L}(C,\sigma)$, which is defined in (3.3).
{toappendix}
Proof of Theorem 4.3.
Suppose $J>1$ steps of the algorithm are performed and consider $2\leq j\leq J$. Let $\Upsilon_{j+1}$ be the function added to $\Upsilon$ in step $j$. During step $j-1$, we start with some $\Upsilon_{j-1}$ and then form a $\frac{d}{2^{j-1}(C+1)}$-packing $M_{j-1}$ of the set $B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$. We then set $\Upsilon_{j}=\mathop{\mathrm{argmin}}_{f\in M_{j-1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$.
Applying Lemma 2.7 with $\delta=\frac{d}{2^{j-1}(C+1)}$, $\mathcal{F^{\prime}}:=B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$, and conditioning on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq d/2^{j-2}$ (which implies $\bar{f}\in\mathcal{F}^{\prime}$), we obtain
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}},\Upsilon_{j-1}\right)$ $\leq 3\cdot|M_{j-1}|\exp(-n\varepsilon_{j}^{2})$
where we define
$\varepsilon_{j}$ $=\frac{d\sqrt{L(C,\sigma,B_{\mathcal{F}})}}{2^{j-1}(C+1)}.$
Further, define $a_{j}=\exp(-n\varepsilon_{j}^{2})$.
Now, we upper bound $|M_{j-1}|$ using Lemma 4.2. In the notation of the lemma, we would set $\varepsilon=\frac{d}{2^{j-2}}$, $\delta=\frac{d}{2^{j-1}}$, $f=\Upsilon_{j-1}$, and $g=\bar{f}$. Then $M_{j-1}$ corresponds to a $\varepsilon/c$-packing of $B(f,\varepsilon)$, and we further condition on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}=\|g-f\|_{L_{2}(\mathbb{P}_{X})}\leq\delta$, noting that $\delta<\varepsilon$. Thus Lemma 4.2 implies
$|M_{j-1}|$ $\leq M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{j-3},4(C+1)\right)$
$\leq M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right).$
The second inequality uses Lemma 2.9 since $j\leq J$. The resulting probability upper bound does not depend on $\Upsilon_{j-1}$, so we have proven that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right)\exp(-n\varepsilon_{j}^{2}).$
If we set $A_{j}$ to be the event $\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}$, then we can use the property that $P(A_{J})\leq P(A_{J}|A_{J-1}^{c})+P(A_{j-1}|A_{j-2}^{c})+\dots+P(A_{3}|A_{2}^{c})+P(A_{2})$ to conclude that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right)\cdot\sum_{j=2}^{J}a_{j}$
$\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right)\frac{a_{J}}{1-a_{J}}.$
The final inequality used Lemma 2.10. Note that $a_{J}<1$ since $C$ is large enough that $L(C,\sigma,B_{\mathcal{F}})\neq 0$.
Note that
$\frac{d}{2^{J-3}}=\frac{4\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}.$
Now, if $n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{4\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},4(C+1)\right)$ and $a_{J}=\exp\left(-n\varepsilon_{J}^{2}\right)<\frac{1}{2}$, which implies $(1-a_{J})^{-1}<2$, then we have
@@
$M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{d}{2^{J-3}},4(C+1)\right)\cdot\frac{a_{J}}{1-a_{J}}$ $=M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{4\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},4(C+1)\right)\cdot\frac{a_{J}}{1-a_{J}}$
$<\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\frac{a_{J}}{1-a_{J}}$
$\leq\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\exp\left(-n\varepsilon_{J}^{2}\right)\cdot\frac{1}{1-a_{J}}$
$\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Since $a_{J}<\frac{1}{2}$ iff $n\varepsilon_{J}^{2}>\log 2$, we conclude that if
@@
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{4\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},4(C+1)\right)\vee\log 2,$

### v5

@@ Appendix C Proofs for Section 4
$\leq 3\cdot\left[\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{d}{2^{j-2}},4c)\right]^{2}\exp(-\Lambda(C,\sigma,L)n\delta^{2})$
@@
$=3\cdot\left[\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{c\varepsilon_{j}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},4c)\right]^{2}\exp(-\Lambda(\tfrac{c}{2}-1,\sigma,L)n\delta^{2})$
using non-increasing properties from Lemma 4.3.
For $j=2$, recall $\mathbb{P}(A_{2}\cap A_{1}^{c})=\mathbb{P}(A_{2})$ since $\mathbb{P}(A_{1})=0$. Well $\Upsilon_{2}=f_{i^{\ast}}$ was picked from a maximal $d/c$-packing (hence covering) of $B(f_{R},d)\cap\mathcal{F}$, without pruning. But $B(f_{R},d)\cap\mathcal{F}=B(\bar{f},d)\cap\mathcal{F}$, so the cardinality is bounded by $\mathcal{M}^{\operatorname{adloc}}(\bar{f},d,c)\leq\mathcal{M}^{\operatorname{adloc}}(\bar{f},d,4c)$. Hence applying Lemma 3.9 with $\delta=\tfrac{d}{c}=\tfrac{d}{2(C+1)}$,
$\mathbb{P}(A_{2})$ $\leq\mathbb{P}(\|\bar{f}-f_{i^{\ast}}\|_{L_{2}(\mathbb{P}_{X})}>d/2)$
$=\mathbb{P}(\|\bar{f}-f_{i^{\ast}}\|_{L_{2}(\mathbb{P}_{X})}>(C+1)\delta)$
$\leq 3|\mathcal{O}(f_{R})|\exp(-\Lambda(C,\sigma,L)n\delta^{2})$
$\leq 3\cdot\mathcal{M}^{\operatorname{adloc}}(\bar{f},d,4c)\exp(-\Lambda(C,\sigma,L)n\delta^{2})$
@@
$\leq 3\cdot\left[\mathcal{M}^{\operatorname{adloc}}(\bar{f},d,4c)\right]^{2}\exp(-\Lambda(C,\sigma,L)n\delta^{2})$
@@
$=3\cdot\left[\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{c\varepsilon_{2}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},4c)\right]^{2}\exp(-n\varepsilon_{2}^{2}).$
Part 2: Bounding $\mathbb{P}(A_{J})$ for $1\leq J\leq J^{\ast}$.
Set $a_{J}=\exp(-n\varepsilon_{J}^{2})$. Summing over $2\leq j\leq J$ and using Lemma 4.3 to replace $\varepsilon_{j}$ with $\varepsilon_{J}$ in the local metric entropy, we conclude as in the non-adaptive case that
@@
$\mathbb{P}(A_{J})$ $\leq\sum_{j=2}^{J}\mathbb{P}(A_{j}\cap A_{j-1}^{c})\leq 3\sum_{j=2}^{J}\left[\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{c\varepsilon_{j}}{\sqrt{\Lambda(C,\sigma,L)}},4c)\right]^{2}\exp(-n\varepsilon_{j}^{2})$